A train runs at a speed of 10 km per hour greater than half the speed of the other train. The former runs 175 km in 5 hours. What is the speed of the latter.
Answer by checkley77(7072)

175/5=35 kmh. for the former train.
10+35/2=10+17.5=27.5 kmh. for the latter train.